Jet’s Pizza is considering investing in a new type of pizza oven. If the discount rate is 7% for these mutually exclusive projects which project should the company invest in?

Modern Brick Oven Brick Oven
Time 0 -10,000 -10,000
Time 1 5,000 4,000
Time 2 4,000 3,000
Time 3 3,000 10,000

Question 15 options:

a)

Modern Oven

b)

Neither project — both have negative NPV.

c)

Both projects — both have positive NPV.

d)

Brick Oven
Answer #1

F726 fic =SUM(F722:F725) B D 712 713 714 PV@7% PVCF $ (10,000.00) 715 716 717 718 Modern Brick Oven Year Cashflows 0 $ (10,00F726 for =SUM(F722:F725) 712 713 714 Modern Brick Oven Year PV@7% 715 -1 Cashflows -10000 5000 4000 717 718 =E715/1.07 =E716/